The following table shows the current status data for a project at the end of month 3.

PV

(Planned Value)

EV

(Earned Value)

AC

(Actual Cost)

SV

CV

SPI

CPI

$18,461.54

$35,673.08

$20,673.08

If you were the project manager for this project, should you have any concerns about this data?

If yes, then

  • what are your concerns?
  • what might be some underlying causes?
  • how might you combat these in the future?

If no, then explain why not.
Answer #1

Earned Value (EV) = 35,673.08

Planned Value (PV) = 18,461.54

Actual Cost (AC) = 20,673.08

a) Schedule Variance = Earned Value to date - Planned Value to date

If SV > 0, the week's activity is ahead of the schedule

If SV < 0, the week's activity is behind the schedule

If SV = 0, the week's activity is on the schedule

Here, Schedule Variance = EV - PV = 35,673.08 - 18,461.54

Schedule Variance = $17211.54

Hence, the project is ahead of the schedule.

b) Cost Variance = Earned Value to date - Actual Cost to date

If CV > 0, the week's activity is under budget

If CV < 0, the week's activity is over budget

If CV = 1, the week's activity is on budget

Here, Cost Variance = EV - AV = 35,673.08 - 20,673.08

Cost Variance = $15000

Hence, the project is under budget.

c) Schedule Performance Index = Earned Value to date / Planned Value to date

If SPI > 1, the week's activity is ahead of the schedule

If SPI < 1, the week's activity is behind the schedule

If SPI = 1, the week's activity is on the schedule

Here, Schedule Performance Index = EV / PV = 35,673.08 / 18,461.54

Schedule Performance Index = 1.932

Hence, the project is ahead of the schedule.

d) Cost Performance Index = Earned Value to date / Actual Cost to date

If CPI > 1, the week's activity is under budget

If CPI < 1, the week's activity is over budget

If CPI = 1, the week's activity is on budget

Here, Cost Performance Index = EV / AC = 35,673.08 / 20,673.08

Cost Performance Index = 1.726

Hence, the project is under budget.

Because the project is well ahead of the schedule and project is under budget, the project manager should be relieved given the project is over-achieving the target.

Here, the assumption is that the Planned Value, Earned Value and Actual Cost are captured accurately,

The only thing that should concern the manager is that maybe the project budgeting was not appropriate as the earned value is almost double the planned value and hence, they might have to revise the budget otherwise the variance values will be too large at the end of the project.

Kindly up-vote this answer if it was helpful to you. This will help me in maintaining the quality of my answers. Do reach out to me in case of any queries. Thank you.

Happy Learning!